A 0.632 gram of a metal. M. (Momie mass = 154.2) i
Answer #1

1)

m = 0.632 g of metal

m = 0.923 g of MCl

a)

%metal = m/total M = 0.632/0.923 *100 = 68.4723%

2)

empirical formula

mol of M = mass/MW = 0.632/154.2 = 0.004098

mol of Cl = mass/MW

mass of Cl = 0.923 -0.632 = 0.291

mol of Cl = mass/MW = 0.291/35.5 = 0.008197

then

ratios

0.008197/0.004098 = 2

then

MCl2 is the empirical formula